OpenStudy (anonymous):

Evaluate the function rule for the given value y = 4 x 2^x for x= -6

OpenStudy (jhannybean):

\(\large y=4(2^x)\)?

OpenStudy (jhannybean):

Is that the problem?

OpenStudy (anonymous):

y= 4 times 2 to the power of x

OpenStudy (anonymous):

out of these: 1/8 1/32 -48 1/16 by "/" i mean a fraction

OpenStudy (jhannybean):

ok, so \[\large y=4(2^{(-6)})\]\[\large y=4*\frac{1}{2^{6}}\]\[\large 2^{6}= 64\]\[\large y=4*\frac{1}{64}\]simplify and you'll have your answer

OpenStudy (jhannybean):

you can rewrite \[\large 2^{-6} \to \frac{1}{2^6}\]

OpenStudy (anonymous):

1/8?

OpenStudy (jhannybean):

how many times does 4 go into 64?

OpenStudy (anonymous):